### v1

# Learning to Explore and Select for Coverage-Conditioned Retrieval-Augmented Generation

Abstract
Interactions with billion-scale large language models typically yield long-form responses due to their extensive parametric capacities, along with retrieval-augmented features. While detailed responses provide insightful viewpoint of a specific subject, they frequently generate redundant and less engaging content that does not meet user interests. In this work, we focus on the role of query outlining (i.e., selected sequence of queries) in scenarios that users request a specific range of information, namely coverage-conditioned (C2) scenarios. For simulating C2 scenarios, we construct QTree, 10K sets of information-seeking queries decomposed with various perspectives on certain topics. By utilizing QTree, we train QPlanner, a 7B language model generating customized query outlines that follow coverage-conditioned queries. We analyze the effectiveness of generated outlines through automatic and human evaluation, targeting on retrieval-augmented generation (RAG). Moreover, the experimental results demonstrate that QPlanner with alignment training can further provide outlines satisfying diverse user interests. Our resources are available at https://github.com/youngerous/qtree.
Learning to Explore and Select for Coverage-Conditioned
Retrieval-Augmented Generation
Takyoung Kim1,† Kyungjae Lee2 Young Rok Jang2 Ji Yong Cho2,3Gangwoo Kim4,†Minseok Cho2Moontae Lee2,5 1University of Illinois Urbana-Champaign 2LG AI Research 3Cornell University 4Korea University 5University of Illinois Chicago tk30@illinois.edu   moontae.lee@lgresearch.ai
Work done as a research intern at LG AI Research.
1 Introduction
Curiosity is attributed to perceptual knowledge and sequence of interests (BERLYNE, 1954). The emerging performance of recent large language models (LLMs) enables providing long and detailed responses by utilizing their deep parametric knowledge. Meanwhile, human-machine interaction interface (e.g., chat and acoustic interface) that has been studied for a long time (Levin et al., 2000) allows users to request customized information in the form of instructions. More recent proprietary chat interface services such as ChatGPT (OpenAI, 2023), Gemini (Google, 2024), and BingChat are accelerating the exploration of human curiosity.
LLMs’ long responses consist of innumerable combinations of relevant topics, as there is no gold standard in long-form text (Krishna et al., 2021; Xu et al., 2022, 2023). However, LLMs often fail to provide tailored responses when users’ queries get complicated. For instance, a user can request LLMs to provide information on a topic, Generative AI, especially focusing on historical context without describing recent trends (since the user already understands the latest news), but it is too complicated for LLMs to generate meticulously-composed responses. Since humans have their own background knowledge and curiosity level, we define these queries that constrain the information coverage of certain topics as coverage-conditioned (C2) queries. Specifically, our scenario conceptualizes “coverage” by assuming that users can request LLMs to include or exclude a specific subtopic within long-form responses.
Figure 1: Unlike the existing RAG systems that simply rewrite or decompose the given query, our QPlanner, trained with QTree, first explores possible information coverage. A subset of queries is selected as an outline to retrieve relevant passages and generate long-form response. By leveraging retrieved passages, we prompt LLM to generate the final response tailored to user interests.
There have been several studies to generate responses following users’ instructions (Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022a; Taori et al., 2023; Asai et al., 2023). Among these trends, we focus on query outlining, which is a sequence of queries representing intermediate subtopics within a long text, demonstrating its effectiveness in long-form story generation (Fan et al., 2018; Sun et al., 2022; Yang et al., 2022, 2023; Wang et al., 2024; Shao et al., 2024; Lee et al., 2024). On top of that, we further raise a question: “How to generate customized outlines for long-form responses in the retrieval-augmented setup?” Well-generated outlines should align with user perspectives (Shao et al., 2024) and show improved performance in downstream tasks, such as document retrieval and response generation. In our work, outlines consisting of relevant topics to C2 queries are expected to be utilized both as better search queries and intermediate content drafts, thereby improving the end-to-end retrieval-augmented generation (RAG) performance.
To explore the way to generate customized outlines, we propose QTree, 10K hierarchical sets of information-seeking queries interpreting user queries of existing datasets with diverse perspectives. Hierarchies of QTree are comprised based on the abstraction level of the main topic. For example, as illustrated in Figure 1, hierarchical topics related to processes after pretraining (i.e., Finetuning and RLHF) are selected among various viewpoints in LLMs. We assume that the selected subset of QTree performs a role as an outline for document search and response generation.
To extract a subset of queries following C2 query from QTree, we construct QPlanner, an on-premise autoregressive 7B language model generating customized and preferable outlines without human intervention. Specifically, QPlanner generates outlines to arbitrary C2 queries by learning the outline-extracting capabilities of LLMs. We evaluate the performance of QPlanner by measuring whether generated outlines follow C2 queries, both with automatic and human evaluation. Moreover, we verify that training QPlanner with preference alignment can further improve the outline quality and end-to-end RAG output. Our contributions are summarized as follows:
  • We explore the effectiveness of query outlining in coverage-conditioned (C2) scenarios by constructing QTree, 10K hierarchical sets of queries containing various perspectives.
  • We train QPlanner, an autoregressive 7B language model that generates outlines for long-form responses customized to C2 queries.
  • By conducting automatic and human evaluation, we verify the effectiveness of alignment training for outlining C2 queries and analyze how the aligned outlines help generate long-form responses, focusing on the RAG task.
2 Related Work
2.1 Query Modification with LLMs
Integrating retrieval systems with LLMs has become crucial in enhancing retrieval-augmented systems, where query reformulation plays a pivotal role in improving information retrieval outcomes. Recent advancements involve prompting LLMs to supply detailed information, such as expected documents or pseudo-answers, for query expansion (Wang et al., 2023; Jagerman et al., 2023). Additionally, there’s a growing interest in utilizing reward signals to fine-tune the query reformulation model, further optimizing search results based on the ranking of retrieved documents (Ma et al., 2023; Yoon et al., 2024). Recent efforts have also focused on breaking down complex questions into simpler subqueries, leveraging the decomposition for more accurate retrieval and response generation (Khot et al., 2023). Additionally, a line of studies (Yao et al., 2023; Press et al., 2023; Khattab et al., 2023) merge the concept of Chain-of-Thoughts (CoT) (Wei et al., 2022b) with interaction through retrieval systems. Furthermore, most recent studies (Oh et al., 2024; Weller et al., 2024) introduce the inclusion of instructional scenarios in the information-seeking process.
Figure 2: The overview of our framework. [Step 1] Base query ($q_{base}$) is decomposed into queries with diverse viewpoints (QTree), preceded by generating coverage query ($q_{cov}$). [Step 2] After C2 candidate outlines are extracted, an LLM evaluates each outline and selects the best-scored one. [Step 3] Utilizing this dataset, QPlanner is trained to sequentially generate its own QTree and preferred outline by taking the C2 query as an input.
While having similar query modification strategies (e.g., query decomposition or instructional scenarios), we consider query outlining as an effective strategy before generating long-form responses. Previously, a hierarchical structure has shown remarkable performance on story generation task (Fan et al., 2018). Yang et al. (2022, 2023) have also emphasized the importance of outline for structured story generation. More recently, Shao et al. (2024) have explored the role of the outline as a medium to present diverse perspectives and constructed outlines by iterative conversational processes. Furthermore, Lee et al. (2024) have improved the quality of free-form writing by outline augmentation. Above these studies, we position our work as the automatic exploration of a better outline for retrieval-augmented long-form generation in C2 scenarios.
2.2 Evaluation of Long-form Responses
Evaluating long-form responses of LLMs presents a significant challenge due to the subjective and complex nature of such assessments. Recent research (Krishna et al., 2021; Xu et al., 2022, 2023) highlight the limitations of current automated metrics in accurately assessing long-form texts, underscoring the necessity for detailed, fine-grained evaluation methods. Some holistic approaches consider a wide array of model capabilities, introducing multi-metric evaluation settings (Liang et al., 2023; Gehrmann et al., 2023; Shevlane et al., 2023; Ye et al., 2024). Additionally, the push towards fine-grained atomic evaluation settings for tasks like fact verification and summarization indicates a growing interest in detailed assessments (Min et al., 2023; Krishna et al., 2023). Recent research has also explored model-based evaluation, where learned models determine automated scores (Yuan et al., 2021; Liu et al., 2023; Kim et al., 2024). While these studies consider long-form responses themselves as the target to evaluate, we evaluate both outlines and responses by following insights from Kellogg (1988) where outlines alleviate the cognitive overload of reading long text (i.e., serving as an effective content draft) while representing the core structure of long-form writing.
3 Framework
3.1 Definition
Throughout our work, we call QTree as a hierarchical set of queries derived from a single user query. Formally, QTree consists of a hierarchical query sequence $\left\{q_{1},q_{2},...,q_{11},...,q_{ijk}\right\}$, where the length of each query annotation denotes the hierarchical level within QTree. For example, $q_{213}$ refers to the query located at the third position at depth-3, following the path from the second position at depth-1 to the first position at depth-2. We set the depth and the width of QTree as three, respectively (i.e., 3+9+27=39 queries exist in each QTree). Additionally, as exemplified in Figure 1, we define C2 query as the concatenation of user’s original query (base query; $q_{base}$) and additional query constraining information coverage (coverage query; $q_{cov}$), formalized as $C^{2}=[q_{base};q_{cov}]$.
Dataset Source Train Test
ASQA (Stelmakh et al., 2022) Wikipedia 4,353 100
Longform (Köksal et al., 2024) Wikipedia 4,483 100
ExpertQA (Malaviya et al., 2023) Expert 1,741 100
Total - 10,577 300
Table 1: Basic statistics of our seed datasets. We specify the number of questions in each split. We obtain $q_{base}$ from these datasets constructed from various corpus.
3.2 Overview
Figure 2 illustrates the overall pipeline of our framework. Followed by $q_{base}$ collection (Section 3.3.1), we construct QTree (Section 3.3.2) and generate $q_{cov}$ (Section 3.3.3). Since we intend to generate C2 queries dealing with users’ various requests, QTree has to interpret $q_{base}$ with diverse perspectives. For example, a $q_{base}$What is Generative AI?" can be answered through various viewpoints, including its latest trend, historical context, and application across different fields. By utilizing generated $q_{cov}$, candidate outlines for each C2 query can be parsed (Section 3.4.1), preceded by the evaluation of each outline for selecting the best one (Section 3.4.2). The following subsections detail the construction procedure, and all used prompts are provided in Appendix E.
3.3 Preparing C2 Queries (Step 1)
3.3.1 Base Query ($q_{base}$) Collection
We first collect $q_{base}$ that requires long-form content composition to respond. Specifically, we utilize two Wikipedia-based long-form question answering datasets – ASQA (Stelmakh et al., 2022) and Longform (Köksal et al., 2024), and one from expert domain – ExpertQA (Malaviya et al., 2023), as demonstrated in Table 1. For the test set, we sample 100 test queries for each dataset. By leveraging LLMs [^1] We fix the LLM with the gpt-4-0125-preview of OpenAI (OpenAI, 2023), with a temperature of 1.0, throughout this work., we construct C2 queries by combining these 10K $q_{base}$ with corresponding $q_{cov}$ (see Section 3.3.3). Since there are a few noisy queries, we slightly modify and filter existing queries. We describe the detailed preprocessing in Appendix A.
3.3.2 QTree Construction
Prior to generating $q_{cov}$, we decompose $q_{base}$ into diverse queries as a tree structure (i.e., QTree). The purpose of constructing QTree for each $q_{base}$ is to unfold the scope of information within parametric knowledge of LLMs. This structured graph also enables effective instruction generation (will be detailed in the next subsection) according to the hierarchy of abstractiveness. We assume queries in deeper depth present more concrete subtopics. In other words, queries in leaf nodes contain the most detailed queries of a specific subject. Table 6 in Appendix C illustrates an example of QTree.
Quality Check
Checking the quality of QTree is relatively easier than unformatted text sequence. In the query decomposition stage, we ensure that QTree contains a predefined number of nodes in each depth and width and does not overlap each other. This can be simply done by heuristically inspecting the structured output. We also prompt the LLM to generate queries as question-format [^2] We intend every query contains question mark, but we do not explicitly filter queries with non-question format as questions can be expressed in various forms..
3.3.3 Coverage Query ($q_{cov}$) Generation
Intent Operation $q_{cov}$ Examples
[…]
      any thematic discussions by experts on the qualifications or
      contributions of the newly appointed UPSC member to the commission.
[…]
      the name Jibril originated from.
[…]
      seasons can enhance or diminish the quality and quantity of tea
      produced in various regions.
[…]
      only interested in the birth date of the actress who played the bird lady.
[…]
      to focus on the different classifications of theft in various legal systems
      without delving into the proof elements.
[…]
      I’m only interested in the one who directed the initial movie.
Table 2: Example of generated $q_{cov}$ according to intent operations (randomly sampled from the training set).
C2 Query: Describe the film The Woman Hunt. Since I’m already familiar with how audiences and critics received The Woman Hunt, please avoid discussing reviews or reception in your explanation.
Parsed Outline: 1. What is the plot of The Woman Hunt? 1.1. What are the main events in The Woman Hunt? 1.1.1. What initiates the conflict in The Woman Hunt? 1.1.2. What is the climax of The Woman Hunt?
Table 3: Example of parsed outline. Example of corresponding QTree is available at Table 6.
To remind, $q_{cov}$ constrain coverage within the wide range of information. Therefore, generating $q_{cov}$ from QTree requires selecting specific viewpoint to cover. However, solely relying on LLMs’ parametric knowledge does not guarantee the diversity of realistic situations. We therefore adopt the following two concepts to assist $q_{cov}$ generation.
  • Background Query: Understandably, users’ interests attribute to their background knowledge. With this in consideration, we randomly select a single query from QTree, which will be a starting point for simulating users’ curiosity. We define this query as background query. The specificity of the background query also differs according to the depth of the selected query. In other words, it can be inferred that users who are seeking information of deeper queries possess a more detailed recognition of the relevant subjects.
  • Intent Operation: While users consider a particular subject to ask, they may choose whether the content should be addressed within the responses. We conceptualize user intent through a binary operation (i.e., Inclusion, Exclusion), thereby facilitating the generation of $q_{cov}$ that explicitly request the inclusion/exclusion of the subtopic on the background query.
In practice, we prompt LLM to generate $q_{cov}$ by combining randomly selected background query from QTree with intent operation. As shown in Table 2, combinations of background query and intent operation yield diverse $q_{cov}$ for each $q_{base}$. Especially, in Exclusion operation, we find that relatively harder requirements (e.g., avoiding one topic but focusing on another topic) are given. Therefore, it can be assumed that $q_{cov}$ with Exclusion operation requires harder scenarios (we analyze this in Appendix D). In addition, we generate five $q_{cov}$ per each $q_{base}$ and choose one if corresponding outlines are correctly parsed (will be further described in Section 3.4.1).
3.4 Exploring Candidate Outlines & Evaluation (Step 2)
3.4.1 Parsing Outlines
In this stage, LLM sequentially extracts candidate outlines from QTree that satisfies instructions of C2 queries [^3] Our preliminary verification identifies that sequentially generating candidate outlines shows more diversity than temperature-based sampling. Refer to Appendix B for case studies.. Table 3 illustrates the example of the candidate outline, providing a hierarchical subset of queries (i.e., plot - main event - conflict & climax) about The Woman Hunt. We set the number of candidate outlines to three per each C2 query.
Quality Check
We fix the number of queries within each outline to four, guaranteeing that all queries are directly connected or neighboring within QTree, as illustrated in Table 3. Additionally, we verify the JSON parsability of each outline and ensure that all queries do not overlap each other. For the efficient usage of API calls, we heuristically remove queries in leaf nodes if an outline contains more than four queries and include the outline as a candidate.
3.4.2 Evaluating Outline Quality
In order to evaluate the quality of candidate outlines, we leverage LLM (gpt-4-0125-preview) to serve as a judge deciding whether the content on each candidate outline follows C2 query. More precisely, we prompt the model to assign scores, measuring how faithfully the outline aligns with the C2 query. The scoring mechanism employs a five-point Likert scale, along with a corresponding rationale generation for the evaluation. Since outlines have significantly shorter context than long-form text while containing core contents (Kellogg, 1988), it is expected that the evaluation of the LLM shows better alignment with human intuition than directly evaluating long context.
3.5 Training QPlanner (Step 3)
To deal with arbitrary C2 queries and generate corresponding outlines, we train a language model named QPlanner. We instruct QPlanner to sequentially generate QTree and selected outline, as we intend that QTree serves like an intermediate Chain-of-Thought (Wei et al., 2022b) reasoning process.
In the training phase, we employ supervised fine-tuning (SFT) and alignment tuning for the training QPlanner. First, we train the Llama-2-7B-Chat model (Touvron et al., 2023) using 10K C2 queries mapped with 31K candidate outline pairs, which are constructed through Section 3.4.1. This training phase allows the model to generate formatted outlines following C2 queries. Then we further align the preferred outline by adopting a variant of direct preference optimization (DPO) (Rafailov et al., 2023). Following Tunstall et al. (2023) that show the possibility of distilling the preference of large teacher models into the targeted model, we utilize LLM evaluation previously acquired in Section 3.4.2 as a reward signal for aligning QPlanner. We regard the chosen sample as the highest-scored and the rejected sample as the lowest-scored outline. We do not use samples whose highest and lowest scores are the same in the alignment stage. The amount of the final training sample is 31,488 in SFT and 8,568 in DPO, respectively. Refer to Appendix F for further details, such as hyperparameters.
4 Experiments
Mean ($\uparrow$) SD ($\downarrow$)
Random Basline 2.57 1.44
SFT-QPlanner 2.79 1.40
DPO-SynNeg 2.98 1.39
(31K + 8K align)
DPO-Combined 3.01 1.36
(31K + 16K align)
DPO-QPlanner 3.16 1.33
(Ours; 31K + 8K align)
Table 4: Mean and standard deviation (SD) for automatic outline evaluation (five-point Likert scale). Our DPO-QPlanner scores the highest mean score and the lowest SD, indicating robust improvement.
4.1 Automatic Evaluation
We let LLM (gpt-4-0125-preview) conduct automatic scoring for outlines on our test set. For long-form response evaluation, we do not automatically measure due to the lack of reliability in long-form text evaluation (Xu et al., 2023).
4.1.1 Mean Score Comparison
Table 4 shows our test result with a five-point Likert scale. We score outlines generated by each trained model [^4] A few cases return outline with 3 or 5 queries, which is not an ideal number of the output (i.e., 4), but we do not filter them in our evaluation., focusing on whether the content of outlines follows given C2 queries (as mentioned in Section 3.4.2).
DPO-QPlanner (Ours)
While our SFT model shows a higher score than the random baseline, we find that preference alignment based on the SFT model using our candidate outlines, which are ranked in Section 3.5, significantly improves the evaluation score. It should be noted that our reward signal is based on the language model’s generation whose subjectivity is similar to human’s, different from Ma et al. (2023); Yoon et al. (2024) whose reward signals are more likely to be an explicit and gold criterion (  e.g., retrieved ranking of the target document). Since there is no proper baseline to verify the effectiveness of our outlining framework, we construct the following baselines as an initial point to estimate performance.
Random Baseline
Since the output of QPlanner includes QTree, we can extract arbitrary outline regardless of the C2 queries. Specifically, we select a random background query from QTree generated by our SFT-QPlanner, then extend the branch to randomized directions (i.e., upper depth, neighbor, or lower depth) until four queries are connected as a single outline [^5] Intent operation is also not considered in the random baseline.. We find that the random baseline shows the lowest mean score (2.57) with the highest standard deviation (1.44) on our test set.
DPO-SynNeg
To further explore the effectiveness of selected negative samples in DPO-QPlanner, we prepare another DPO model trained with different types of negative samples. While negative samples of DPO-QPlanner are generated by LLM, as described in Section 3.5, it is also possible to heuristically synthesize negative samples using QTree, background query, and intent operation. This procedure is similar to generating random baseline, except for (1) setting the starting point as the background query and (2) ensuring synthesized outlines have the opposite intent operation to the original intent. That is, the intent of synthetic negative (SynNeg) outlines explicitly conflicts with that of the positive (chosen) outlines, while the intent of negative outlines of DPO-QPlanner are aligned with that of positive outlines. On the 8K DPO-QPlanner training set, we maintain the chosen (i.e., highest-scored) outlines and replace rejected outlines with synthetically generated negative outlines.
On the comparison with DPO-QPlanner and DPO-SynNeg, we find that negative samples of DPO-QPlanner are notably effective than the other one. We conjecture that constructing “hard” negatives [^6] We define LLM-generated negative samples as “hard” when compared to synthetic negative samples whose intent operations are explicitly opposite to positive outlines. (i.e., less scored subtrees with the “same” intent) is an important factor to align with C2 queries, as discussed in previous studies (Rosset et al., 2023; Scarlatos et al., 2024).
DPO-Combined
We also measure the performance of combining negative samples in DPO-QPlanner and DPO-SynNeg. While the test score slightly increases compared to DPO-SynNeg, performance becomes worse than DPO-QPlanner despite the doubled amount of training samples.
4.1.2 Pairwise Score Comparison
We also measure pairwise comparison between models in Figure 3. Our DPO-QPlanner gets the highest score on the same C2 query, which is intuitively aligned with the result in Table 4. Given the fact that exhaustively devising preferred outline for C2 queries is labor-intensive, our QPlanner is an effective solution for exploring and providing preferred outlines for composing long-form responses.
Figure 3: Pairwise score comparison per each C2 query in automatic outline evaluation. Outlines of our DPO-QPlanner are judged as the best among all other baselines.
4.2 Human Evaluation
We also conduct a human study to identify the effectiveness of QPlanner. We describe detailed information in Appendix G such as demographics and payment.
4.2.1 Outline Evaluation
Recruited evaluators read and rate two outlines (SFT- and DPO-QPlanner) for randomly selected 100 C2 queries from the test set using the same five-point Likert scale criteria as the automatic evaluation. Each evaluator rates from 4 to 10 outlines in random order, and each outline has at least 6 evaluators (Average: 6.55, Max: 15). We intend to have as many evaluators as possible to rate individual outline to gather a collective rating for each set. This is because even with simple outlines, judging if outlines with unfamiliar subjects successfully follow C2 queries is highly intellectual and unavoidably subjective task.
As expected for this type of subjective task, evaluators do not reach a high inter-rater reliability (Krippendorff’s $\alpha$=0.22 for SFT ratings 0.23 for DPO-QPlanner ratings between human raters). However, more importantly, we find significant positive correlations between human-rated scores and LLM-rated scores – both for SFT (Pearson’s $r$ = 0.51, p-value $<$ 0.001) and DPO-QPlanner (Pearson’s $r$ = 0.39, p-value $<$ 0.001), which indicates positive relationships with large and medium strength, respectively. Moreover, DPO-QPlanner receives higher scores (Mean= 3.29, Std=0.81) than SFT (Mean= 3.03, Std=0.78). This trend is supported even when we regress scores on model version (SFT or DPO-QPlanner) and the total length of outlines in characters, with outline id as a fixed effect (Model:$b$=0.27, p-value=0.01; Outline length: $b$=0.01, p-value=0.30). That is, the length of outlines is not predictive of scores.
4.2.2 Response Evaluation
In addition to outline evaluation, we recruit another participants to evaluate 100 randomly sampled RAG responses. Ten evaluators are assigned for each response. According to Kim et al. (2023), the writing format of model responses affects human preferences. Therefore, we fix the format with Markdown to compare responses by focusing only on their content. In addition, we prompt LLM to generate responses by strictly relying on given evidence to prevent LLM from arbitrarily responding with its parametric knowledge. We provide a description of the document retrieval procedure in Section G.2.1.
Regarding evaluation criteria, we first instruct participants to judge whether generated responses follow requirements within C2 queries or not (Query Satisfaction in Figure 4). We guide to annotate "Yes" if responses at least partially handle topics within C2 queries. For response pairs annotated as "Yes" in both models, participants select their preferred response from the two models (Response Preference in Figure 4).
QPlanner as Better Search Query
We verify whether queries within outlines can help search relevant documents. We compare responses of vanilla RAG with those of QPlanner-DPO using the exactly same prompt. In other words, outline queries of QPlanner-DPO only affect search results. As illustrated in Figure 4(a), we find that conventional RAG pipeline does not properly handle C2 queries, whose requirements are far more complicated than normal queries. Furthermore, among responses that satisfy the requirements of C2 queries, responses of QPlanner-DPO are mostly preferred.
QPlanner as Better Content Draft
We further explore whether better outlines lead to better responses. In this setup, we compare responses of QPlanner-SFT with those of QPlanner-DPO. Likewise, the exactly same prompt is used for the comparison and queries within outlines are included to compose responses, along with document searching. Results in Figure 4(b) indicate that QPlanner aligned with preference outlines tends to provide preferred outlines, although QPlanner-SFT also shows a similar tendency with QPlanner-DPO, which is different from that of Vanilla-RAG.
(a) QPlanner as Search Query
(b) QPlanner as Content Draft
Figure 4: Human evaluation results
For both studies, we conduct two McNemar’s tests (McNemar, 1947) to examine whether the differences we found are statistically significant. We choose to run McNemar’s tests because they are suited for our experiment settings, where the outcome is binary (i.e., satisfactory or not), and each human evaluator judges both responses for the same query (within-subjects design). The contingency tables used for the tests can be found in Table 10. The results confirm that QPlanner-DPO outperforms Vanilla RAG (test-statistics=$60$, p-value<0.001), and QPlanner-DPO outperforms QPlanner-SFT (test-statistics=$207$, p-value=0.004).
5 Conclusion
Customizing information coverage within C2 queries requires complex selection of subtopics to compose long-form responses. We investigate the role of query outline in terms of assistive search queries and intermediate content draft for RAG task. To simulate C2 scenarios, we construct QTree, hierarchical sets of queries representing diverse perspectives of the original query. Subsequently, we train QPlanner which extracts customized outlines from QTree for C2 queries. Our findings based on automatic and human evaluation show that (1) preference-aligned QPlanner can provide better outlines, (2) outlines enable searching more relevant documents, and (3) well-generated outlines lead to preferred responses. We believe our work shows the possibility of QTree as a testbed for exploring effective pre-writing strategy to deal with complicated queries.
Limitation
We fix the number of queries in each outline to four, but it can be adjusted (i.e., composing more or less queries) according to the topic. We plan to adapt outlines in a flexible manner in our future work.
It should be noted that retrieved documents in our RAG setup can also affect final responses. Although we set the same searching configuration among all methodologies, additional fact verification is still needed for practical applications.
In addition, despite demonstrating significant performance gain both in automatic and human evaluation, we find that state-of-the-art LLM still has difficulty in generating long-form responses that handle detailed coverage of C2 query. This is presumably due to the complexness of C2 queries, but it arises the importance of constructing meticulous benchmark evaluating long-form responses of complicated queries. This will be another direction of our future work.
Appendix A Base Query Modification
For the ASQA [^7] Apache 2.0 License (Stelmakh et al., 2022) and ExpertQA [^8] MIT License (Malaviya et al., 2023) dataset, we do not modify the base query. For the Longform [^9] MIT License (Köksal et al., 2024) dataset, as there are additional format-related instructions concatenated with the base query (e.g., Respond in 3 sentences.), we eliminate them by using regular expressions. Moreover, we find that Longform dataset contains noisy queries (e.g., This does not provide enough information for an answer to be provided.), which are unfiltered artifacts generated by large language models. In this case, we manually filter similar expressions.
Appendix B Outline Sampling Comparison
To identify the effectiveness of sequentially generating candidate outlines at once, we generate candidate outlines using temperature sampling. As shown in Table 5, sequential generation meaningfully diversifies candidate outlines than temperature sampling. Similar results are obtained when we increase the sampling temperature.
C2 Query: When was the museum of islamic art built? Since I’m already familiar with who the lead architect of the Museum of Islamic Art was and their design vision, focus on providing the construction date without delving into architectural design specifics.
Parsed Outline (Sampling Temperature 1.0): 3.2. Which construction company undertook the building of the Museum of Islamic Art, and in which year did they commence work? 3.2.1. What firm was responsible for the construction of the Museum of Islamic Art, and when did they start? 3.2.2. Can you identify the construction firm that built the Museum of Islamic Art and the year construction began? 3.2.3. Who constructed the Museum of Islamic Art, and in what year did construction efforts begin? 3.3. What were the financial and logistical challenges in constructing the Museum of Islamic Art? 3.3.1. What budgetary constraints affected the construction of the Museum of Islamic Art? 3.3.2. What logistical hurdles were encountered during the building of the Museum of Islamic Art? 3.3.3. In constructing the Museum of Islamic Art, what were the major financial and logistical issues 3. Who contributed to the design and construction of the Museum of Islamic Art? 3.2. Which construction company undertook the building of the Museum of Islamic Art, and in which year did they commence work? 3.2.1. What firm was responsible for the construction of the Museum of Islamic Art, and when did they start? 3.2.2. Can you identify the construction firm that built the Museum of Islamic Art and the year construction began?
Candidate Outlines (Sequential Generation): 2. When did the Museum of Islamic Art open to the public? 2.1. What is the official opening date of the Museum of Islamic Art? 2.1.1. On which date was the Museum of Islamic Art inaugurated? 2.1.2. When was the grand opening of the Museum of Islamic Art? 1. What is the construction date of the Museum of Islamic Art? 1.2. Who was the architect behind the Museum of Islamic Art design, and in what year did the construction take place under their guidance? 1.2.1. What is the name of the architect who designed the Museum of Islamic Art, and during which year was their design realized? 1.2.3. In what year did construction of the Museum of Islamic Art occur under the designated architects design? 3. Who contributed to the design and construction of the Museum of Islamic Art? 3.2. Which construction company undertook the building of the Museum of Islamic Art, and in which year did they commence work? 3.2.1. What firm was responsible for the construction of the Museum of Islamic Art, and when did they start? 3.2.3. Who constructed the Museum of Islamic Art, and in what year did construction efforts begin?
Table 5: Comparison of temperature sampling and sequential generation of candidate outlines.
Appendix C Example of QTree
C2 Query: Describe the film The Woman Hunt. Since I’m already familiar with how audiences and critics received The Woman Hunt, please avoid discussing reviews or reception in your explanation.
QTree: 1. What is the plot of The Woman Hunt? 1.1. What are the main events in The Woman Hunt? 1.1.1. What initiates the conflict in The Woman Hunt? 1.1.2. What is the climax of The Woman Hunt? 1.1.3. How does The Woman Hunt end? 1.2. Who are the main characters in The Woman Hunt? 1.2.1. Who is the protagonist of The Woman Hunt? 1.2.2. Who is the antagonist in The Woman Hunt? 1.2.3. What supporting characters play crucial roles in The Woman Hunt? 1.3. What themes are explored in The Woman Hunt? 1.3.1. What is the primary theme of The Woman Hunt? 1.3.2. How does The Woman Hunt explore gender dynamics? 1.3.3. What messages does The Woman Hunt convey about survival? 2. Who directed The Woman Hunt? 2.1. What is the directorial style of The Woman Hunt? 2.1.1. How does the director use camera angles in The Woman Hunt? 2.1.2. What unique directorial choices are made in The Woman Hunt? 2.1.3. How does the pace affect the narrative in The Woman Hunt? 2.2. What other films has the director of The Woman Hunt made? 2.2.1. What are the most popular films by The Woman Hunts director? 2.2.2. How do other films by the director compare to The Woman Hunt? 2.2.3. What recurring themes appear in the directors filmography? 2.3. How has the directors background influenced The Woman Hunt? 2.3.1. What aspects of the directors personal life reflect in The Woman Hunt? 2.3.2. How does the directors cultural background inform The Woman Hunt? 2.3.3. What previous experiences of the director shaped The Woman Hunt? 3. How was The Woman Hunt received by audiences and critics? 3.1. What are the critical reviews of The Woman Hunt? 3.1.1. How do film critics analyze The Woman Hunt? 3.1.2. What are the predominant critiques of The Woman Hunt? 3.1.3. Are there any notable defenses of The Woman Hunts thematic choices? 3.2. What is the audiences reaction to The Woman Hunt? 3.2.1. How do audience perspectives on The Woman Hunt vary? 3.2.2. What aspects of The Woman Hunt resonate most with audiences? 3.2.3. What fan opinions of The Woman Hunt diverge from critical reviews? 3.3. Has The Woman Hunt won any awards or recognition? 3.3.1. What awards or nominations has The Woman Hunt received? 3.3.2. How does The Woman Hunt rank among other films of its genre? 3.3.3. Are there any film festivals where The Woman Hunt was highlighted?
Table 6: Example of QTree generated by the process described in Section 3.3.2.
Appendix D Analysis on Intent Operations
      SFT       DPO-QPlanner
      Inclusion       Exclusion       Inclusion       Exclusion
      Mean       2.85       2.74       3.22       3.10
      SD       1.23       1.55       1.15       1.47
Table 7: Mean and standard deviation (SD) according to the intent operation in automatic outline evaluation.
We decompose the result of Table 4 according to intent operations in Table 7, focusing on SFT and DPO-QPlanner scores. We discover that C2 queries based on Exclusion score lower than those on the intent of Inclusion. This result aligns with our assumption in Section 3.3.3, where $q_{cov}$ with Exclusion operation require more complicated selection of desirable outline.
Appendix E Used Prompts
We use the following prompts in our work.
Appendix F Training Details
      Hyperparameter       SFT       DPO
      Epoch       1       1
      Batch Size Per Device       14       8
      Learning Rate (LR)       2e-5       5e-7
      LR Schedule       Cosine       Cosine
      Warmup Ratio       0.1       0.1
      Gradient Accumulation Step       1       2
      Beta       -       0.01
      # of Samples       31,488       8,568
Table 8: Hyperparameters for training QPlanner. A few noisy samples are filtered in advance at SFT stage.
We utilize publicly available software [^10] https://github.com/huggingface/alignment-handbook in our finetuning and alignment stage. We train each stage one epoch using 16 NVIDIA A100 GPUs (40GB of memory). Table 8 describes detailed hyperparameters for each stage.
Appendix G Additional Information on Human Evaluation
We conduct two human evaluation studies separately. For both studies, crowd workers are recruited from Prolific [^11] https://www.prolific.com/. At the beginning of the evaluation, workers are informed what task they are expected to do, there are no foreseeable benefits and risks, their participation is voluntary, and they can leave if they want (see Figure 5).
G.1 Outline Evaluation
Score Rubric
1: The sub-questions/responses entirely disregards the instructions, providing content unrelated to the instruction. 2: The sub-questions/responses show a superficial attempt to follow instructions but significantly strays from the intended task, missing key objectives. 3: The sub-questions/responses generally adheres to the instructions but overlooks certain details or nuances, achieving only a partial match with the instruction. 4: The sub-questions/responses is closely aligned with the instructions, exhibiting minor deviations that slightly affect the completeness of the execution. 5: The sub-questions/responses exhibits impeccable adherence to the instructions, capturing all nuances and completing the task as specified.
Table 9: Score rubric for evaluating subqueries and responses in human evaluation.
A total of 127 crowd workers participate in the evaluation (Gender: 68 men, 57 women, and 2 non-binary; Age: Mean=28.6 yrs, SD=7.9 yrs, Min=18 yrs, Max=63 yrs; Ethnicity: White: 69, Black: 44, Mixed: 11, and Asian:2; Country of residence: South Africa: 52 (41.27%), Portugal: 20 (15.87%), Poland: 10 (7.94%), United Kingdom: 5 (3.97%), Mexico: 5 (3.97%), and 19 other countries; Highest education level completed: A majority of the evaluators hold at least a Bachelor’s degree (n=83, 65.87%)). Individual crowd workers evaluate different numbers of instances depending on their availability. They are compensated 9 GBP/hour for their work. We paid 606.41 GBP in total.
They score each sample by following the rubric in Table 9. We engage evaluators by asking them to write at least 30 characters to describe their rationale for preference, which also helps evaluators take the rating more seriously and derive more rational and accurate ratings. We also provide an evaluation session with no more than evaluating 5 outlines considering the human attention span. If evaluators want to continue participating, they have to sign up for another evaluation session, which ensures they have a break for recharging themselves. Lastly, we utilize Prolific’s offering that automatically rejects work that takes too long or too short, above or below two standard deviations of the average completion time.
G.2 Response Evaluation
Search Query Evaluation Content Draft Evaluation
QPlanner-DPO QPlanner-DPO
Vanilla RAG Unsatisfactory Satisfactory QPlanner-SFT Unsatisfactory Satisfactory
Unsatisfactory 179 606 Unsatisfactory 130 271
Satisfactory 60 155 Satisfactory 207 392
  • Note: The sum of the counts in the contingency tables is 1000 (100 queries evaluated by 10 workers) for each evaluation, respectively.
Table 10: Contingency tables for human study in response evaluation
A total of 63 crowd workers participate in the evaluation (Gender: 33 men and 30 women; Age: Mean=28.08 yrs, SD=9.01 yrs, Min=19 yrs, Max=68 yrs; Ethnicity: White: 26, Black: 22, Mixed: 10, and Asian: 5; Country of residence: South Africa: 20 (15.87%), Portugal: 10 (7.94%), Mexico: 8 (6.35%), Poland: 6 (4.76%), Canada: 4 (3.17%), and 11 other countries; Highest education level completed: A majority of the evaluators hold at least a Bachelor’s degree (n=49, 77.78%)). Individual crowd workers evaluate different numbers of instances depending on their availability. They are compensated 9 GBP/hour for their work. We pay 878.58 GBP in total.
We offer ten single sessions for evaluation (5 sessions for Vanilla RAG vs. QPlanner-DPO and 5 sessions for QPlanner-SFT vs. QPlanner-DPO). Each session has ten evaluators. If wanted, evaluators can participate in more than one session; thirteen out of 63 evaluated multiple sessions. A session takes 20 to 40 minutes. In a session, evaluators are first provided with a short tutorial with evaluation guidelines and examples. Then, they evaluate a pair of responses that answer the same query for twenty queries. They are told that all of the queries are formatted as [question] ($q_{base}$) + [instruction] ($q_{cov}$) and asked to mark a response “Satisfactory” if the response satisfies any of the two evaluation items: (1) the response indeed answers the question [question] with evidence or partial evidence, which includes “there is no evidence but here is useful information,” and (2) the response follows the instruction, [instruction], and mark it “Unsatisfactory” otherwise. If all of the two responses to the same query are rated “Satisfactory,” they were asked to choose which one was a better answer (155 cases in Search Query Evaluation and 392 cases in Response Outline Evaluation; refer to Table 10). We describe a better answer would have more items described above satisfied, or it would be better at following the instruction.
G.2.1 Document Retrieval
For simulating RAG pipeline, we utilize DuckDuckgo [^12] https://serpapi.com/duckduckgo-search-api to search relevant documents. To balance the number of documents, top-10 documents are retrieved in Vanilla RAG, and top-2 documents are retrieved for queries (i.e., 2 * 5 = 10 documents including C2 query) in QPlanner-SFT and QPlanner-DPO. Additionally, we follow the associative selection process, suggested in Lee et al. (2023), to extract relevant evidence paragraph from retrieved documents. Specifically, we construct FLAN-T5-Large (Chung et al., 2024) trained with Wikipedia-based datasets such as MS-MARCO (Bajaj et al., 2018), ELI5 (Fan et al., 2019), ASQA (Stelmakh et al., 2022), and Qasper (Dasigi et al., 2021). The trained language model matches passages in each document with given queries and returns answerability score deciding whether the paired query and passage are relevant. We select top-1 passage for each document as an evidence for generating RAG response.
Figure 5: Initial information provided to participants in our human study.