### v2

# Learning to Explore and Select for Coverage-Conditioned Retrieval-Augmented Generation

Abstract
Interactions with large language models (LLMs) often yield long and detailed responses, leveraging both parametric knowledge and retrieval-augmented generation (RAG). While these responses can provide rich insights, they often include redundant or less engaging content not aligned with user interests. This issue becomes apparent when users specify particular subtopics to include or exclude – termed coverage-conditioned (C2) queries – as LLMs often struggle to provide tailored responses. To address this challenge, we investigate the role of query outlines, sequences of subqueries designed to guide LLMs in generating responses that meet specific user requirements. To systematically create and evaluate these outlines, we introduce QTree, a dataset of 10K hierarchical sets of information-seeking subqueries that define structured boundaries for outline creation and evaluation in C2 scenarios [^1] Our resources are available at https://github.com/youngerous/qtree.. Additionally, we develop QPlanner, a 7B language model trained to generate customized outlines within boundaries of QTree. We evaluate the effectiveness of the generated outlines through automatic and human judgements, focusing on their impact within retrieval-augmented generation (RAG) systems. Experimental results demonstrate that QPlanner, especially when trained with alignment techniques like DPO, generates higher-quality outlines that better fulfill diverse user needs.
Learning to Explore and Select for Coverage-Conditioned Retrieval-Augmented Generation
Takyoung Kim1,† Kyungjae Lee2 Young Rok Jang2 Ji Yong Cho2,3Gangwoo Kim4,†Minseok Cho2Moontae Lee2,5 1University of Illinois Urbana-Champaign 2LG AI Research 3Cornell University 4Korea University 5University of Illinois Chicago tk30@illinois.edu   moontae.lee@lgresearch.ai
Work done as a research intern at LG AI Research.
1 Introduction
Figure 1: QTree constrains the range of available outlines for the user’s C2 query, and tailored outlines satisfying the requirement of C2 query are selected for RAG downstream tasks.
Recent advancements of large language models (LLMs) have enabled them to provide long and detailed responses by leveraging their parametric knowledge. As these models improve, human-machine interaction interfaces (e.g., chat and acoustic interfaces) – which have been studied for a long time (Levin et al., 2000) – have become more sophisticated, allowing users to request highly specific and personalized information. Proprietary chat services such as ChatGPT (OpenAI, 2023), Gemini (Google, 2024), and BingChat have further accelerated the exploration of personalized information. Additionally, retrieval-augmented generation (RAG) methods are being adopted to enhance the relevance and timeliness of LLM responses by integrating external knowledge.
Despite these advancements, LLMs often struggle with delivering tailored responses when faced with complex user queries. For instance, a user might request LLMs to provide information on Generative AI, focusing specifically on its historical context while excluding recent trends. Crafting such meticulously composed responses is difficult for LLMs for two reasons: (1) LLMs’ long-form outputs can contain innumerable combinations of relevant topics, and (2) there is no established gold standard for long-form text generation (Krishna et al., 2021; Xu et al., 2022, 2023). Recognizing this, we first define queries that constrain the information coverage of certain topics as coverage-conditioned (C2) queries, where “coverage” refers to the user’s intent to instruct LLMs to include or exclude specific subtopics within their responses. These C2 queries especially pose challenges in constructing long-form RAG responses as they require selective document retrieval as well.
To improve LLM responses for users’ complex queries, there have been works on query outlining, creating sequences of intermediate subtopics to guide long-form responses. Query outlining has been effective in areas like long story generation (Fan et al., 2018; Sun et al., 2022; Yang et al., 2022, 2023; Wang et al., 2024; Shao et al., 2024). However, generating high-quality outlines that address complex queries like C2 queries remains challenging, as there is no systematic approach for creating and evaluating such outlines.
With the concepts of C2 query and query outlining in place, we pose two key research questions:
  1. RQ1.
    How can we create and evaluate better outlines for C2 queries?
  2. RQ2.
    Can these outlines improve RAG systems by serving as search queries and content drafts?
To address RQ1, we present QTree, a dataset comprising 10K hierarchical sets of information-seeking subqueries (with 39 subqueries in each set) that interpret user queries with diverse perspectives, facilitating the exploration and selection of appropriate outlines for C2 queries. The hierarchies in QTree are organized according to the abstraction level of the main topic, defining tangible boundaries of available outlines. For example, as illustrated in Figure 1, hierarchical subtopics related to processes after pretraining (i.e., Fine-tuning and RLHF) are selected as proper outlines for RAG response among various viewpoints on the topic of Training LLMs, following the requirements of the C2 query. In contrast, less relevant subtopics in QTree (e.g., Pretraining LLMs) will not be a desirable outline for the C2 query. By leveraging QTree, we can systematically create and judge outlines for long-form responses, ensuring that they align with the user’s coverage constraints.
Regarding RQ2, we introduce QPlanner, an autoregressive 7B language model designed to generate tailored outlines within QTree’s hierarchical boundaries. We hypothesize that high-quality outlines aligned with C2 queries can improve both document retrieval and response generation in RAG systems. We also evaluate QPlanner ’s performance through both automatic metrics and human judgments, assessing the quality of the generated outlines and their impact on downstream tasks. Experimental results on C2 queries from diverse domains (i.e., Wikipedia and expert domains) demonstrate that training QPlanner with preference alignment further improves both outline quality and overall RAG performance.
Our contributions are summarized as follows:
  1. 1.
    We present QTree, a novel dataset of 10K hierarchical subquery sets that define boundaries for available outlines, facilitating the creation and evaluation of better outlines for coverage-conditioned (C2) queries (addressing RQ1).
  2. 2.
    We introduce QPlanner, an autoregressive language model designed to generate customized outlines that improve document retrieval and content generation in RAG systems (addressing RQ2).
  3. 3.
    We conduct comprehensive evaluations, including automatic metrics and human judgments, to validate the effectiveness of our approach in enhancing outline quality and RAG performance.
2 Related Work
Figure 2: The overview of our framework. [Step 1] Base query ($q_{base}$) is decomposed into subqueries with diverse viewpoints (QTree), preceded by generating coverage query ($q_{cov}$). [Step 2] After C2 candidate outlines are extracted, a judge LLM evaluates each outline and selects the best-scored one. [Step 3] Utilizing this dataset, QPlanner is trained to sequentially generate its own QTree and preferred outline by taking the C2 query as an input.
2.1 Query Modification with LLMs
Integrating retrieval systems with LLMs has become crucial, with query modification playing a pivotal role in improving information retrieval outcomes. Recent advancements focus on prompting LLMs to provide detailed information, such as expected documents or pseudo-answers, for query expansion (Wang et al., 2023; Jagerman et al., 2023). Furthermore, reward signals are being used to fine-tune query modification models, optimizing search results based on the ranking of retrieved documents (Ma et al., 2023; Yoon et al., 2024). Additionally, complex questions are being decomposed into simpler subqueries to enhance retrieval accuracy and response generation (Khot et al., 2023).
Among various query modification strategies, query outlining stands out as an effective method for generating long responses. Outlining techniques have been primarily studied in tasks such as story generation (Fan et al., 2018). Yang et al. (2022, 2023) have also emphasized the importance of outline for narrative generation, while Shao et al. (2024) explored the outline as tools for presenting diverse perspectives through iterative conversational processes. More recently, Lee et al. (2024) improved free-form writing with outline augmentation. However, despite these advances, it has seen less attention in retrieval-augmented contexts. In addition, none of these studies systematically evaluate the generated outlines in complex scenarios (e.g., C2 scenarios). Our work aims to address this gap by proposing a controlled evaluation testbed for outlines and their impact on long-form responses.
2.2 Evaluation of Long-form Responses
Evaluating long-form responses from LLMs presents a significant challenge due to the subjective and multifaceted nature of the task. Previous studies (Krishna et al., 2021; Xu et al., 2022, 2023) highlight the limitations of automated metrics in accurately assessing long-form texts, underscoring the need for more nuanced evaluation methods. Several approaches have emerged to tackle this issue by incorporating multi-metric evaluation frameworks (Liang et al., 2023; Gehrmann et al., 2023; Shevlane et al., 2023; Ye et al., 2024), as well as task-specific metrics for fact verification and summarization (Min et al., 2023; Krishna et al., 2023). Recent research has also investigated model-based evaluations where learned models are used to generate automated scores (Yuan et al., 2021; Liu et al., 2023; Kim et al., 2024a).
While most of these studies focus solely on the evaluation of long-form responses, we extend this by evaluating both the outlines and responses they guide. Drawing from insights in cognitive psychology (Kellogg, 1988), we argue that outlines alleviate the cognitive overload for readers, functioning as effective content drafts and providing the core structure for long-form writing.
3 Framework
3.1 Background
We refer to QTree as a tree-shaped hierarchical set of subqueries (defining “subquery” as each node in QTree) derived from a single user query. We set both the depth and the width of QTree at three levels (i.e., 3+9+27=39 subqueries in each QTree). Additionally, as illustrated in Figure 1, we define C2 query as the concatenation of the user’s original query (base query; $q_{base}$) and additional coverage-constraining query (coverage query; $q_{cov}$), represented as $C^{2}=[q_{base};q_{cov}]$.
Dataset Source Train Test
ASQA (Stelmakh et al., 2022) Wikipedia 4,353 100
Longform (Köksal et al., 2024) Wikipedia 4,483 100
ExpertQA (Malaviya et al., 2024) Expert 1,741 100
Total - 10,577 300
Table 1: Basic statistics of our seed datasets. We specify the number of questions in each split. We obtain $q_{base}$ from these datasets constructed from various corpus.
3.2 Overview
Figure 2 illustrates the procedural framework, including the construction of QTree and QPlanner. Followed by collecting $q_{base}$ (Section 3.3.1), we construct QTree (Section 3.3.2) and generate $q_{cov}$ (Section 3.3.3). Generated C2 queries (i.e., $q_{base}$ and $q_{cov}$) are then utilized to select candidate outlines. For example, the answer to a $q_{base}$What is Generative AI?" can contain diverse perspectives, including its latest trend, historical context, and application across different fields. Within available outlines that guide to satisfying answers, our goal is to obtain the outline that follows $q_{cov}$Tell me about its historical context". Therefore, within the range of QTree, we parse candidate outlines for each C2 query (Section 3.4.1), preceded by the evaluation for selecting the optimal outline (Section 3.4.2). The following subsections detail the procedural generation, and all used prompts are provided in Appendix E.
3.3 Preparing C2 Queries (Step 1)
3.3.1 Base Query ($q_{base}$) Collection
We first collect $q_{base}$ that requires long-form content composition to respond. Specifically, we employ two Wikipedia-based long-form question answering datasets – ASQA (Stelmakh et al., 2022) and Longform (Köksal et al., 2024), and one from expert domains – ExpertQA (Malaviya et al., 2024), as demonstrated in Table 1. For the test set, we sample 100 test queries for each dataset. By leveraging LLMs [^2] We use gpt-4-0125-preview of OpenAI (OpenAI, 2023) with a temperature of 1.0, throughout this work., we construct C2 queries by combining these 10K $q_{base}$ with corresponding $q_{cov}$ (see Section 3.3.3). We slightly modify and filter a few $q_{base}$ containing noises, described in Appendix A.
3.3.2 QTree Construction
Prior to generating $q_{cov}$, we decompose $q_{base}$ into diverse subqueries as a tree structure (i.e., QTree). The purpose of constructing QTree for each $q_{base}$ is to unfold the scope of information within parametric knowledge of LLMs. This structured graph also enables effective instruction generation (will be detailed in the next subsection) according to the hierarchy of abstractiveness. Subqueries in deeper depth present more specific subtopics. Table 6 in Appendix C illustrates an example of QTree.
Quality Check
In the query decomposition stage, we ensure that QTree contains a predefined number of subqueries (i.e., three) in each depth and width and does not overlap each other. This can be simply done by heuristically inspecting and comparing the structured output.
3.3.3 Coverage Query ($q_{cov}$) Generation
Intent Operation $q_{cov}$ Examples
[…]
      any thematic discussions by experts on the qualifications or
      contributions of the newly appointed UPSC member to the commission.
[…]
      the name Jibril originated from.
[…]
      seasons can enhance or diminish the quality and quantity of tea
      produced in various regions.
[…]
      only interested in the birth date of the actress who played the bird lady.
[…]
      to focus on the different classifications of theft in various legal systems
      without delving into the proof elements.
[…]
      I’m only interested in the one who directed the initial movie.
Table 2: Example of generated $q_{cov}$ according to intent operations (randomly sampled from the training set).
C2 Query: Describe the film The Woman Hunt. Since I’m already familiar with how audiences and critics received The Woman Hunt, please avoid discussing reviews or reception in your explanation.
Parsed Outline: 1. What is the plot of The Woman Hunt? 1.1. What are the main events in The Woman Hunt? 1.1.1. What initiates the conflict in The Woman Hunt? 1.1.2. What is the climax of The Woman Hunt?
Table 3: Example of parsed outline. Example of corresponding QTree is available at Table 6.
To remind, the role of $q_{cov}$ is to specify certain subtopics to address (i.e., include or exclude) within a broad range of information. Therefore, generating $q_{cov}$ from QTree requires selecting a specific viewpoint to cover. However, solely relying on LLMs’ parametric knowledge does not guarantee the diversity of realistic situations. We therefore adopt the following two concepts to assist in generating $q_{cov}$.
  • Background Subquery: Understandably, asking for specific knowledge means that users are recognizing the knowledge itself. With this in consideration, we randomly select a single subquery from QTree, which will be the knowledge users are aware of. We define this subquery as background subquery. The specificity of the background subquery differs according to the depth of the selected query.
  • Intent Operation: While considering a particular subject to ask, users may choose whether the content should be addressed within the responses. We conceptualize user intent through a binary operation (i.e., Inclusion, Exclusion), thereby facilitating the generation of $q_{cov}$ that explicitly request the inclusion/exclusion of the subtopic on the background subquery.
In practice, we prompt LLM to generate $q_{cov}$ by combining a randomly selected background subquery from QTree with intent operation [^3] Although we use background subquery to generate $q_{cov}$ in this section, it is also used to construct baselines. Refer to Section 4.2.. As demonstrated in Table 2, combinations of background subquery and intent operation yield diverse $q_{cov}$ for each $q_{base}$. Especially, requirements of $q_{cov}$ with Exclusion operation are more complicated (e.g., avoiding one topic but focusing on another topic) than Inclusion. We analyze the performance difference according to intent operations in Appendix D. We sample five preliminary $q_{cov}$ per each $q_{base}$ and finally choose one if corresponding three candidate outlines are parsed correctly (which will be further described in Section 3.4.1).
3.4 Exploring Candidate Outlines & Evaluation (Step 2)
3.4.1 Parsing Outlines
In this stage, LLM sequentially extracts JSON-formatted candidate outlines from QTree that satisfy instructions of C2 queries [^4] Our preliminary verification identifies that sequentially generating candidate outlines shows more diversity than temperature-based sampling. Refer to Appendix B for case studies.. Table 3 visualizes the example of a candidate outline, consisting of hierarchical subqueries (i.e., plot - main event - conflict & climax) about The Woman Hunt. We extract three candidate outlines per each C2 query.
Quality Check
We fix the number of subqueries within each outline to four, guaranteeing that all subqueries are directly connected or neighboring within QTree, as illustrated in Table 3. Additionally, we verify the JSON parsability of each outline and ensure that all subqueries do not overlap each other. For the efficient usage of API calls, we heuristically remove subqueries in leaf nodes if an outline contains more than four subqueries and include the outline as a candidate.
3.4.2 Evaluating Outline Quality
In order to rank three candidate outlines, we leverage LLM (gpt-4-0125-preview) to serve as a judge deciding whether the content on each candidate outline follows C2 query. More precisely, we prompt the model to assign five-point Likert-scale scores with rationales, measuring how faithfully the outline aligns with the C2 query. Since outlines are significantly shorter than long-form text while maintaining core contents (Kellogg, 1988), it is expected that evaluating outlines is more efficient and intuitive than directly evaluating long responses. These scored outlines are utilized as supervision and alignment pairs for training QPlanner, which will be described in further sessions.
3.5 Training QPlanner (Step 3)
To generalize with arbitrary C2 queries, we train a 7B language model named QPlanner. We instruct QPlanner to sequentially generate QTree and select an outline, as we intend that QTree serves like an intermediate Chain-of-Thought (Wei et al., 2022) reasoning process. More technical details are described in Section 4.1.
4 Experiments
Mean ($\uparrow$) SD ($\downarrow$)
Random Basline 2.57 1.44
SFT-QPlanner 2.79 1.40
DPO-SynNeg 2.98 1.39
(31K + 8K align)
DPO-Combined 3.01 1.36
(31K + 16K align)
DPO-QPlanner 3.16 1.33
(Ours; 31K + 8K align)
Table 4: Mean and standard deviation (SD) for automatic outline evaluation (five-point Likert scale). DPO-QPlanner scores the highest mean score and the lowest SD, indicating robust improvement.
4.1 Training Details of QPlanner
We employ supervised fine-tuning (SFT) and alignment tuning for the training QPlanner. First, we train the Llama-2-7B-Chat model (Touvron et al., 2023) using 10K C2 queries mapped with 31K candidate outline pairs, constructed through Section 3.4.1. This training phase allows the model to generate formatted outlines following C2 queries (named SFT-QPlanner hereafter).
Then we further align the preferred outline by adopting a variant of direct preference optimization (DPO) (Rafailov et al., 2023). Following Tunstall et al. (2024) that show the possibility of distilling the preference of large teacher models into a targeted model, we utilize LLM evaluation scores previously acquired in Section 3.4.2 as reward signals for aligning QPlanner (named DPO-QPlanner hereafter). We regard the highest-scored outline as a positive (chosen) sample and the lowest-scored outline as a negative (rejected) sample. We skip samples whose highest and lowest scores are the same in the alignment stage.
The amount of the final training sample is 31,488 for SFT-QPlanner and 8,568 for DPO-QPlanner, respectively. Refer to Appendix F for further details, such as hyperparameters.
4.2 Baselines for Outline Comparison
Random Baseline
Since the output of QPlanner accompanies QTree as an intermediate reasoning process, we can extract an arbitrary outline by leveraging this, regardless of the C2 queries. Specifically, we select a random background subquery from QTree generated by SFT-QPlanner, then extend the branch to randomized directions (i.e., upper depth, neighbor, or lower depth) until four subqueries are connected as a single outline. Intent operation is not considered in this random baseline.
DPO-SynNeg
To further explore the effectiveness of selected (i.e., LLM-scored) negative samples in DPO-QPlanner, we prepare another DPO model trained with different types of negative samples. While negative samples of DPO-QPlanner are based on LLM scores, we can also heuristically synthesize negative samples with QTree, background subquery, and intent operation. This procedure is similar to generating random baseline, except for ensuring that synthesized outlines have the opposite intent operation to the original intent. For example, if the positive outline includes background subquery, the synthesized outline is designed to exclude that subquery by selecting another random background subquery within QTree. For the opposite situation, the synthesized outline must contain background subquery of the positive outline. On the 8K DPO-QPlanner training set, we maintain the positive samples and replace negative samples with synthetically generated outlines.
DPO-Combined
We also measure the performance of combining negative samples in DPO-QPlanner and DPO-SynNeg. That is, the number of training samples is doubled.
5 Results
5.1 Automatic Outline Evaluation
We prompt LLM (gpt-4-0125-preview) to score generated outlines in the test set. We use scoring rubric in Table 9 in Section G.1.
5.1.1 Mean Score Comparison
Table 4 shows our test result with a five-point Likert scale. We score outlines generated by each trained model [^5] A few cases return an outline with 3 or 5 queries, which is not an ideal number of the output (i.e., 4), but we do not filter them in our evaluation., focusing on whether the content of outlines follows given C2 queries (as mentioned in Section 3.4.2).
We find that the random baseline shows the lowest mean score (2.57) with the highest standard deviation (1.44) on our test set. While SFT-QPlanner shows a higher score than the random baseline, we find that DPO-QPlanner significantly improves the score (3.16) with the lowest standard deviation (1.33). It implies that we can leverage LLM-generated scores as reward signals in query outlining when constructing positive-negative pairs, even in the absence of explicit and gold reward criteria for their construction (Ma et al., 2023; Yoon et al., 2024).
On the comparison with DPO-QPlanner and DPO-SynNeg, we observe that negative samples of DPO-QPlanner are notably more effective than the other one, since the only difference between them is the type of negative samples. We conjecture that constructing “hard” negatives [^6] We define LLM-generated negative samples as “hard” when compared to synthetic negative samples whose intent operations are explicitly opposite to positive outlines. (i.e., less scored subtrees with the “same” intent) is an important factor to align with C2 queries, sharing insights with different studies on hard negatives (Rosset et al., 2023; Scarlatos et al., 2024). Regarding DPO-Combined, the performance becomes worse than DPO-QPlanner despite the doubled amount of alignment pairs, implying the importance of selective negative samples.
Figure 3: Pairwise comparison for each C2 query in automatic outline evaluation.
5.1.2 Pairwise Comparison
For comprehensive evaluation, we also compare pairwise scores among models. As illustrated in Figure 3, generated outlines of DPO-QPlanner are more preferred than all other baselines on the same C2 query, which is aligned with the atomic scoring result in Table 4. Given the fact that exhaustively devising preferred outlines for C2 queries is labor-intensive, our QPlanner is an effective solution for exploring and creating preferred outlines for long-form responses.
5.2 Human Outline Evaluation
We conduct a human study to identify the effectiveness of QPlanner. We describe detailed experimental setup in Appendix G, such as demographics and payment.
We let participants read and rate two outlines generated from SFT- and DPO-QPlanner for randomly selected 100 C2 queries from the test set using the same five-point Likert scale criteria as the automatic evaluation. Each evaluator rates from 4 to 10 outlines in random order (to avoid position bias), and each outline has at least 6 evaluators (Average: 6.55, Max: 15). We intend to have as many evaluators as possible to rate individual outlines to gather a collective rating for each set. This is because even with simple outlines, judging outlines with unfamiliar topics is a highly intellectual and unavoidably subjective task.
Consequently, we find significant positive correlations between human-rated scores and LLM-rated scores – both for SFT (Pearson’s $r$ = 0.51, p-value $<$ 0.001) and DPO-QPlanner (Pearson’s $r$ = 0.39, p-value $<$ 0.001), which indicates positive relationships with large and medium strength, respectively. Moreover, DPO-QPlanner receives higher human scores (Mean= 3.29, Std=0.81) than SFT (Mean= 3.03, Std=0.78) [^7] This trend is supported even when we regress scores on model version (SFT or DPO-QPlanner) and the total length of outlines in characters, with outline id as a fixed effect (Model:$b$=0.27, p-value=0.01; Outline length: $b$=0.01, p-value=0.30). That is, the length of outlines is not predictive of scores.. We report that evaluating highly subjective tasks may introduce varied ratings among human evaluators despite assigning a large number of evaluators to derive the majority opinion (Krippendorff’s $\alpha$: SFT-QPlanner = 0.22; DPO-QPlanner = 0.23), as observed in other studies (Rottger et al., 2022; Abercrombie et al., 2023).
5.3 Human RAG Evaluation
For long-form response evaluation, we do not automatically measure due to the lack of reliability in long-form text evaluation (Xu et al., 2023). Instead, we recruit another participant to validate the effectiveness of QPlanner on RAG downstream tasks [^8] Details such as recruitment, instructions, and compensation are described in Appendix G.2. 100 RAG responses from the test set are sampled for evaluation, and ten evaluators are assigned for each response. Following insights from Kim et al. (2024b) where the writing format of model responses affects human preferences, we fix the response format with Markdown to compare responses by focusing only on their content. In addition, we prompt LLM999gpt-4-0125-preview is used. to generate responses by strictly relying on given evidence to prevent LLM from arbitrarily responding with its parametric knowledge. We assume web search scenarios for the RAG setup, providing detailed information in Section G.2.1.
Regarding evaluation criteria, we first instruct participants to judge whether generated responses follow requirements of C2 queries or not (Query Satisfaction in Figure 4). We guide them to annotate "Yes" if responses at least partially address topics within C2 queries. For response pairs annotated as "Yes" in both models, participants select their preferred response (Response Preference in Figure 4).
QPlanner as Better Search Query
We verify whether subqueries within outlines can help search relevant documents. We compare responses of vanilla RAG with those of DPO-QPlanner using the exactly same prompt. That is, subqueries of DPO-QPlanner only affect the search result. As illustrated in Figure 4(a), we observe that the conventional RAG pipeline does not properly retrieve relevant evidence for answering C2 queries, whose requirements are far more complicated than normal queries. Furthermore, among responses that satisfy the requirements of C2 queries, responses of DPO-QPlanner are mostly preferred.
QPlanner as Better Content Draft
We further investigate whether better outlines lead to better responses. In this setup, we compare responses of SFT-QPlanner and DPO-QPlanner. The exactly same prompt is used for this comparison, and subqueries within outlines are included in the prompt for composing responses and retrieving documents. Results in Figure 4(b) indicate that further aligning QPlanner with preference can provide preferred outlines, while SFT-QPlanner also shows a similar tendency with DPO-QPlanner.
(a) QPlanner as Search Query
(b) QPlanner as Content Draft
Figure 4: Human evaluation results.
Since the outcome of both studies is binary (i.e., satisfactory or not), and each human evaluator judges both responses for the same query (within-subjects design), we conduct two McNemar’s tests (McNemar, 1947) to examine whether the differences we find are statistically significant. The contingency tables used for the tests can be found in Table 10. The results confirm that DPO-QPlanner significantly outperforms Vanilla RAG (test-statistics = $60$, p-value < 0.001) and SFT-QPlanner (test-statistics = $207$, p-value = 0.004).
6 Conclusion
In this work, we suppose complicated user scenarios asking for a constrained range of a specific topic, called coverage-conditioned (C2) query scenarios. To simulate C2 scenarios and a controlled environment for creating and evaluating query outlines (RQ1), we construct QTree, hierarchical sets of subqueries representing diverse perspectives of the original query. Playing a role as boundaries for available outlines, QTree allows systematic comparison of diverse outlines. Subsequently, we train QPlanner which extracts customized outlines from QTree for C2 queries. Regarding our RQ2, our findings based on automatic and human evaluation show that (1) preference-aligned QPlanner can generate better outlines, (2) outlines enable improved document search, and (3) better outlines lead to preferred responses. We believe our work shows the possibility of QTree as a testbed for exploring effective pre-writing strategies to deal with complicated queries.
Limitations
We discuss the current limitations of our work. First, our graphical representation of subquery nodes adheres to canonical tree structures, with each node connected to three child nodes, but it can be adjusted (i.e., composing more or less subqueries) according to tasks or domains. For example, in a complex domain like medical diagnosis, a larger number of subqueries might be necessary to cover various symptoms, possible conditions, diagnostic tests, and treatment options. In contrast, for a straightforward factual query in a domain like mathematics, fewer subqueries might be sufficient to reach a comprehensive answer. Identifying this optimal number still remains an open question and represents a promising direction for future investigation. We believe that our experimental setup serves as an initial testbed for validating these research questions.
It should also be noted that the contents of retrieved documents in our RAG setup can affect the detailed factual consistency of final responses. Although we set the same search configuration among all methodologies, additional fact verification of documents and responses is still needed for practical applications.
While demonstrating significant performance gains both in automatic and human judgements, we find that state-of-the-art LLMs still have difficulty generating long-form responses that handle detailed coverage of C2 queries. This is presumably due to the complexness of C2 queries, and it arises the importance of constructing meticulous benchmarks evaluating long-form responses to complicated queries. This will be another direction of the future work. Lastly, we would like to mention that our five-point scoring schema can be further improved by considering multiple aspects with a fine-grained score rubric.
Ethical Considerations
Since our QTree is generated with benchmarks based on Wikipedia and domain experts, we do not filter sensitive or unsafe contents throughout our studies. For the practical application in the future, deliberate content selection will be required for the safety. In addition, we explicitly share our experimental setup of human studies for transparency in Appendix G.
Appendix A Base Query Modification
For the ASQA [^10] Apache 2.0 License (Stelmakh et al., 2022) and ExpertQA [^11] MIT License (Malaviya et al., 2024) dataset, we do not modify the base query. For the Longform [^12] MIT License (Köksal et al., 2024) dataset, as there are additional format-related instructions concatenated with the base query (e.g., Respond in 3 sentences.), we eliminate them by using regular expressions. Moreover, we find that Longform dataset contains noisy queries (e.g., This does not provide enough information for an answer to be provided.), which are unfiltered artifacts generated by large language models. In this case, we manually filter similar expressions.
Appendix B Outline Sampling Comparison
To identify the effectiveness of sequentially generating candidate outlines at once, we generate candidate outlines using temperature sampling. As shown in Table 5, sequential generation meaningfully diversifies candidate outlines than temperature sampling. Similar results are obtained when we increase the sampling temperature.
C2 Query: When was the museum of islamic art built? Since I’m already familiar with who the lead architect of the Museum of Islamic Art was and their design vision, focus on providing the construction date without delving into architectural design specifics.
Parsed Outline (Sampling Temperature 1.0): 3.2. Which construction company undertook the building of the Museum of Islamic Art, and in which year did they commence work? 3.2.1. What firm was responsible for the construction of the Museum of Islamic Art, and when did they start? 3.2.2. Can you identify the construction firm that built the Museum of Islamic Art and the year construction began? 3.2.3. Who constructed the Museum of Islamic Art, and in what year did construction efforts begin? 3.3. What were the financial and logistical challenges in constructing the Museum of Islamic Art? 3.3.1. What budgetary constraints affected the construction of the Museum of Islamic Art? 3.3.2. What logistical hurdles were encountered during the building of the Museum of Islamic Art? 3.3.3. In constructing the Museum of Islamic Art, what were the major financial and logistical issues 3. Who contributed to the design and construction of the Museum of Islamic Art? 3.2. Which construction company undertook the building of the Museum of Islamic Art, and in which year did they commence work? 3.2.1. What firm was responsible for the construction of the Museum of Islamic Art, and when did they start? 3.2.2. Can you identify the construction firm that built the Museum of Islamic Art and the year construction began?
Candidate Outlines (Sequential Generation): 2. When did the Museum of Islamic Art open to the public? 2.1. What is the official opening date of the Museum of Islamic Art? 2.1.1. On which date was the Museum of Islamic Art inaugurated? 2.1.2. When was the grand opening of the Museum of Islamic Art? 1. What is the construction date of the Museum of Islamic Art? 1.2. Who was the architect behind the Museum of Islamic Art design, and in what year did the construction take place under their guidance? 1.2.1. What is the name of the architect who designed the Museum of Islamic Art, and during which year was their design realized? 1.2.3. In what year did construction of the Museum of Islamic Art occur under the designated architects design? 3. Who contributed to the design and construction of the Museum of Islamic Art? 3.2. Which construction company undertook the building of the Museum of Islamic Art, and in which year did they commence work? 3.2.1. What firm was responsible for the construction of the Museum of Islamic Art, and when did they start? 3.2.3. Who constructed the Museum of Islamic Art, and in what year did construction efforts begin?
Table 5: Comparison of temperature sampling and sequential generation of candidate outlines.
Appendix C Example of QTree
C2 Query: Describe the film The Woman Hunt. Since I’m already familiar with how audiences and critics received The Woman Hunt, please avoid discussing reviews or reception in your explanation.
QTree: 1. What is the plot of The Woman Hunt? 1.1. What are the main events in The Woman Hunt? 1.1.1. What initiates the conflict in The Woman Hunt? 1.1.2. What is the climax of The Woman Hunt? 1.1.3. How does The Woman Hunt end? 1.2. Who are the main characters in The Woman Hunt? 1.2.1. Who is the protagonist of The Woman Hunt? 1.2.2. Who is the antagonist in The Woman Hunt? 1.2.3. What supporting characters play crucial roles in The Woman Hunt? 1.3. What themes are explored in The Woman Hunt? 1.3.1. What is the primary theme of The Woman Hunt? 1.3.2. How does The Woman Hunt explore gender dynamics? 1.3.3. What messages does The Woman Hunt convey about survival? 2. Who directed The Woman Hunt? 2.1. What is the directorial style of The Woman Hunt? 2.1.1. How does the director use camera angles in The Woman Hunt? 2.1.2. What unique directorial choices are made in The Woman Hunt? 2.1.3. How does the pace affect the narrative in The Woman Hunt? 2.2. What other films has the director of The Woman Hunt made? 2.2.1. What are the most popular films by The Woman Hunts director? 2.2.2. How do other films by the director compare to The Woman Hunt? 2.2.3. What recurring themes appear in the directors filmography? 2.3. How has the directors background influenced The Woman Hunt? 2.3.1. What aspects of the directors personal life reflect in The Woman Hunt? 2.3.2. How does the directors cultural background inform The Woman Hunt? 2.3.3. What previous experiences of the director shaped The Woman Hunt? 3. How was The Woman Hunt received by audiences and critics? 3.1. What are the critical reviews of The Woman Hunt? 3.1.1. How do film critics analyze The Woman Hunt? 3.1.2. What are the predominant critiques of The Woman Hunt? 3.1.3. Are there any notable defenses of The Woman Hunts thematic choices? 3.2. What is the audiences reaction to The Woman Hunt? 3.2.1. How do audience perspectives on The Woman Hunt vary? 3.2.2. What aspects of The Woman Hunt resonate most with audiences? 3.2.3. What fan opinions of The Woman Hunt diverge from critical reviews? 3.3. Has The Woman Hunt won any awards or recognition? 3.3.1. What awards or nominations has The Woman Hunt received? 3.3.2. How does The Woman Hunt rank among other films of its genre? 3.3.3. Are there any film festivals where The Woman Hunt was highlighted?
Table 6: Example of QTree generated by the process described in Section 3.3.2.
Appendix D Analysis on Intent Operations
      SFT-QPlanner       DPO-QPlanner
      Inclusion       Exclusion       Inclusion       Exclusion
      Mean       2.85       2.74       3.22       3.10
      SD       1.23       1.55       1.15       1.47
Table 7: Mean and standard deviation (SD) according to the intent operation in automatic outline evaluation.
We decompose the result of Table 4 according to intent operations in Table 7, focusing on SFT-QPlanner and DPO-QPlanner scores. We discover that C2 queries based on Exclusion score lower than those on the intent of Inclusion. This result aligns with our assumption in Section 3.3.3, where $q_{cov}$ with Exclusion operation require more complicated selection of desirable outline.
Appendix E Used Prompts
We use the following prompts in our work.
Appendix F Training Details
      Hyperparameter       SFT       DPO
      Epoch       1       1
      Batch Size Per Device       14       8
      Learning Rate (LR)       2e-5       5e-7
      LR Schedule       Cosine       Cosine
      Warmup Ratio       0.1       0.1
      Gradient Accumulation Step       1       2
      Beta       -       0.01
      # of Samples       31,488       8,568
Table 8: Hyperparameters for training QPlanner. A few noisy samples are filtered in advance at SFT stage.
We utilize publicly available software [^13] https://github.com/huggingface/alignment-handbook in our finetuning and alignment stage. We train each stage one epoch using 16 NVIDIA A100 GPUs (40GB of memory). Table 8 indicates detailed hyperparameters for each stage.
Appendix G Additional Information on Human Evaluation
We conduct two human evaluation studies separately (i.e., outline and RAG evaluation). For both studies, crowdworkers are recruited from Prolific [^14] https://www.prolific.com/. At the beginning of the evaluation, workers are informed what task they are expected to do, there are no foreseeable benefits and risks, their participation is voluntary, and they can leave if they want (see Figure 5).
G.1 Outline Evaluation
Score Rubric
1: The sub-questions/responses entirely disregards the instructions, providing content unrelated to the instruction. 2: The sub-questions/responses show a superficial attempt to follow instructions but significantly strays from the intended task, missing key objectives. 3: The sub-questions/responses generally adheres to the instructions but overlooks certain details or nuances, achieving only a partial match with the instruction. 4: The sub-questions/responses is closely aligned with the instructions, exhibiting minor deviations that slightly affect the completeness of the execution. 5: The sub-questions/responses exhibits impeccable adherence to the instructions, capturing all nuances and completing the task as specified.
Table 9: Score rubric for evaluating subqueries and responses in human evaluation.
A total of 127 crowd workers participate in the evaluation (Gender: 68 men, 57 women, and 2 non-binary; Age: Mean=28.6 yrs, SD=7.9 yrs, Min=18 yrs, Max=63 yrs; Ethnicity: White: 69, Black: 44, Mixed: 11, and Asian:2; Country of residence: South Africa: 52 (41.27%), Portugal: 20 (15.87%), Poland: 10 (7.94%), United Kingdom: 5 (3.97%), Mexico: 5 (3.97%), and 19 other countries; Highest education level completed: A majority of the evaluators hold at least a Bachelor’s degree (n=83, 65.87%)). Individual crowd workers evaluate different numbers of instances depending on their availability. They are compensated 9 GBP/hour for their work. We paid 606.41 GBP in total.
They score each sample by following the rubric in Table 9. We engage evaluators by asking them to write at least 30 characters to describe their rationale for preference, which also helps evaluators take the rating more seriously and derive more rational and accurate ratings. We also provide an evaluation session with no more than evaluating 5 outlines considering the human attention span. If evaluators want to continue participating, they have to sign up for another evaluation session, which ensures they have a break for recharging themselves. Lastly, we utilize Prolific’s offering that automatically rejects work that takes too long or too short, above or below two standard deviations of the average completion time.
G.2 RAG Response Evaluation
Search Query Evaluation Content Draft Evaluation
DPO-QPlanner DPO-QPlanner
Vanilla RAG Unsatisfactory Satisfactory SFT-QPlanner Unsatisfactory Satisfactory
Unsatisfactory 179 606 Unsatisfactory 130 271
Satisfactory 60 155 Satisfactory 207 392
  • Note: The sum of the counts in the contingency tables is 1000 (100 queries evaluated by 10 workers) for each evaluation, respectively.
Table 10: Contingency tables for human study in response evaluation
A total of 63 crowd workers participate in the evaluation (Gender: 33 men and 30 women; Age: Mean=28.08 yrs, SD=9.01 yrs, Min=19 yrs, Max=68 yrs; Ethnicity: White: 26, Black: 22, Mixed: 10, and Asian: 5; Country of residence: South Africa: 20 (15.87%), Portugal: 10 (7.94%), Mexico: 8 (6.35%), Poland: 6 (4.76%), Canada: 4 (3.17%), and 11 other countries; Highest education level completed: A majority of the evaluators hold at least a Bachelor’s degree (n=49, 77.78%)). Individual crowd workers evaluate different numbers of instances depending on their availability. They are compensated 9 GBP/hour for their work. We pay 878.58 GBP in total.
We offer ten single sessions for evaluation (5 sessions for Vanilla RAG vs. DPO-QPlanner and 5 sessions for SFT-QPlanner vs. DPO-QPlanner). Each session has ten evaluators. If wanted, evaluators can participate in more than one session; thirteen out of 63 evaluated multiple sessions. A session takes 20 to 40 minutes. In a session, evaluators are first provided with a short tutorial with evaluation guidelines and examples. Then, they evaluate a pair of responses that answer the same query for twenty queries. They are told that all of the queries are formatted as [question] ($q_{base}$) + [instruction] ($q_{cov}$) and asked to mark a response “Satisfactory” if the response satisfies any of the two evaluation items: (1) the response indeed answers the question [question] with evidence or partial evidence, which includes “there is no evidence but here is useful information,” and (2) the response follows the instruction, [instruction], and mark it “Unsatisfactory” otherwise. If all of the two responses to the same query are rated “Satisfactory,” they were asked to choose which one was a better answer (155 cases in Search Query Evaluation and 392 cases in Response Outline Evaluation; refer to Table 10). We describe a better answer would have more items described above satisfied, or it would be better at following the instruction.
G.2.1 Document Retrieval
For simulating RAG pipeline, we utilize DuckDuckgo [^15] https://serpapi.com/duckduckgo-search-api to search relevant documents. To balance the number of documents, top-10 documents are retrieved in Vanilla RAG, and top-2 documents are retrieved for each subquery (i.e., 2 * 5 = 10 documents including C2 query) in SFT-QPlanner and DPO-QPlanner. Additionally, we follow the associative selection process, suggested in Lee et al. (2023), to extract relevant evidence paragraphs from retrieved documents. Specifically, we construct FLAN-T5-Large (Chung et al., 2024) trained with Wikipedia-based datasets such as MS-MARCO (Bajaj et al., 2018), ELI5 (Fan et al., 2019), ASQA (Stelmakh et al., 2022), and Qasper (Dasigi et al., 2021). The trained language model matches passages in each document with given subqueries and returns an answerability score deciding whether the paired subquery and passage are relevant. We select the top-1 passage for each document as evidence for generating RAG response.
Figure 5: Initial information provided to participants in our human study.